ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 92BO5 Lara-Aparicio, Miguel, López de Medrano, Santiago and Fuentes-Pardo, Beatriz, Universidad Nacional Autónoma de México, México Some mathematical aspects of the simulation of the motor circadian rhythm in crayfish The motor circadian rhythm in crayfish shows experimentally two stages: a juvenil stage corresponding to a diurnal animal and non-synchronizable by external stimuli, and a mature stage which is nocturnal and synchronizable. Both stages are noisy and show a strong reaction to stimuli in terms of amplitude and level. We construct a mathematical model with two coupled nonlinear oscillators (based on a previous model for the ERG rhythm) which simulates the various aspects of the two stages. The mathematical model is based on a simulation of the Aschoff’s law, and of its possible consequences under periodic stimuli. Partially supported by DGAPA-IN 214697 and Conacyt PC 2153.